# Answer, Assemble, Ace: Understanding How Transformers Answer Multiple Choice Questions

### v1

@@ Abstract
Multiple-choice question answering (MCQA) is a key competence of performant transformer language models that is tested by mainstream benchmarks. However, recent evidence shows that models can have quite a range of performance, particularly when the task format is diversified slightly (such as by shuffling answer choice order). In this work we ask: how do successful models perform formatted MCQA? We employ vocabulary projection and activation patching methods to localize key hidden states that encode relevant information for predicting the correct answer. We find that prediction of a specific answer symbol is causally attributed to a single middle layer, and specifically its multi-head self-attention mechanism. We show that subsequent layers increase the probability of the predicted answer symbol in vocabulary space, and that this probability increase is associated with a sparse set of attention heads with unique roles. We additionally uncover differences in how different models adjust to alternative symbols. Finally, we demonstrate that a synthetic task can disentangle sources of model error to pinpoint when a model has learned formatted MCQA, and show that an inability to separate answer symbol tokens in vocabulary space is a property of models unable to perform formatted MCQA tasks.111Code is available at https://github.com/allenai/understanding_mcqa.
[…]
Answer, Assemble, Ace:
Understanding How Transformers Answer Multiple Choice Questions
Sarah Wiegreffe♡♣ Oyvind Tafjord♡ Yonatan Belinkov♢ Hannaneh Hajishirzi♡♣ Ashish Sabharwal♡ ♡Allen Institute for AI, ♣University of Washington, ♢Technion wiegreffesarah@gmail.com
@@ 1 Introduction
Multiple-choice question answering (MCQA) is a mainstay of language model (LM) evaluation (Liang et al., 2022; Gao et al., 2023; Beeching et al., 2023), not in the least because it avoids the challenges of open-ended text evaluation. One prominent example is the MMLU benchmark Hendrycks et al. (2021), which is considered a meaningful signal for new model releases Lambert (2024a, b).
While early LMs were evaluated on MCQA questions by treating the task as next-token-prediction and scoring each answer choice appended to the input independently, many benchmarks and datasets now evaluate models on their ability to produce the correct answer choice symbol in a single inference run when given the answer choices in the prompt (formatted MCQA; Figs. 1 and 3). Indeed, it has been shown that capable models, and particularly those that have been instruction-tuned, perform well on formatted MCQA tasks Liang et al. (2022); Robinson and Wingate (2023); Wiegreffe et al. (2023). At the same time, despite their strong end-task performance, some models are not robust to seemingly inconsequential format changes, such as shuffling the position of the correct answer choice, changing letters that answers are mapped to, or using different symbols (Pezeshkpour and Hruschka, 2023; Alzahrani et al., 2024; Khatun and Brown, 2024).
@@
In this paper, we set out to interpret internal components to uncover how models promote the prediction of specific specific answer choice symbols. We perform vocabulary projection and activation patching analyses on various model families—Llama 2 (Touvron et al., 2023) and Olmo v1 (Groeneveld et al., 2024) and v1.7 AI2 (2024))—and on three datasets (MMLU Hendrycks et al. (2021), HellaSwag Zellers et al. (2019), and a copying task we create). We find:
@@
Models both encode information needed to predict the correct answer symbol and promote answer symbols in the vocabulary space in a very similar fashion across tasks, even when their overall task performance varies.
@@
Answer symbol production is driven by a sparse portion of the network, dominated by a few attention heads, which specialize in promoting specific answer symbols.
@@
The process for correctly answering more unnatural prompt formats is in some cases more complex – a finetuned Olmo 7B model only begins producing the correct symbols for these prompts in later layers. We observe evidence the model may be making predictions first in the space of expected answer symbols (‘‘A’’, ‘‘B’’) before switching to the symbols provided in the prompt (‘‘C’’, ‘‘D’’).
@@
A simple synthetic task is able to disentangle formatted MCQA performance from performance on a specific dataset, and allows us to narrow down a point during training at which Olmo v1.7 learns formatted MCQA.
5.
@@ 2 Related Work
Behavioral analyses of LM abilities to answer formatted MCQA questions prompt models in a black-box manner by constructing different input prompts and observing how they affect models’ predictions (Wiegreffe et al., 2023; Pezeshkpour and Hruschka, 2023; Sun et al., 2024; Zheng et al., 2024; Alzahrani et al., 2024; Khatun and Brown, 2024; Balepur et al., 2024; Wang et al., 2024). These methods and findings are complementary to ours.
Efforts to interpret transformer LMs for MCQA are limited. Most similar to ours is Lieberum et al. (2023), who study what attention heads at the final token position are doing, and isolate a subset they coin “correct letter heads”. They demonstrate how these heads attend to answer symbol representations at earlier token positions and promote the correct symbol based on its position in the symbol order, though they show this behavior does not hold entirely for answer choices other than {‘‘A’’, ‘‘B’’, ‘‘C’’, ‘‘D’’.} Their experiments are limited to one task (MMLU) and one closed-source model (Chinchilla-70B). Unlike us, they do not inspect model behavior when it is poor, nor do they put forth a hypothesis for model behavior when other answer choice symbols are used. We additionally find diverse roles of important attention heads, with some always promoting specific symbols.
Li and Gao (2024) use vocabulary projection and study the norms of weighted value vectors in multi-layer perceptrons of GPT-2 models in order to localize the production of ‘‘A’’, which GPT-2 models are biased towards producing when prompted 0-shot. They find that updating a single value vector can substantially reduce this label bias. Unlike that work, we investigate how models do symbol binding when they are performant at formatted MCQA; label bias is one reason why a model may not be correctly performing symbol binding.
Outside of MCQA, many works investigate how models internally build up representations that lead to their predictions, such as on factual recall Geva et al. (2021); Dai et al. (2022); Geva et al. (2022b); Meng et al. (2022); Geva et al. (2023); Yu et al. (2023); Merullo et al. (2024b) or basic linguistic Vig et al. (2020); Wang et al. (2023); Merullo et al. (2024a); Prakash et al. (2024); Todd et al. (2024) or arithmetic Stolfo et al. (2023); Hanna et al. (2023); Wu et al. (2023) tasks. MCQA differs because it is both directly represented in real benchmarks and a more challenging task.
@@ 3 Formatted MCQA
We first introduce task notation, datasets, and our protocol for selecting strong models to analyze.
Formatted MCQA is a prompt format in which possible answer choices are presented to the model as an enumerated list, each associated with an answer choice symbol. All of our dataset instances are formatted as follows, unless otherwise noted:
For each of the following phrases, select the best completion (A or B). <in-context examples> Phrase: <question> Choices: A: <correct answer> B: <incorrect answer> The correct answer is: where ‘‘A’’.222The actual token is ‘‘ A’’ in the tokenizers of the models we study, but we leave out preceding spaces for simplicity. We include 3 in-context examples for each instance with labels ‘‘A’’, ‘‘B’’, ‘‘B’’.
Models are evaluated on their ability to assign higher probability to the symbol associated with the correct answer choice (here, ‘‘A’’) as opposed to the alternative symbol (here, ‘‘B’’). We compute as ‘‘A’’ if , else ‘‘B’’.
We convert datasets with answer choices per instance to binary multiple-choice by pairing a randomly-selected incorrect answer choice with the correct answer choice. We restrict instances to two options because the quantity we will plot when measuring discriminativity of model components (Section 6) only takes value for choices, but values for (see App. Section A.1 for an elaboration).
@@ 3.1 Datasets
[…]
@@ 3.2 Models
Olmo.
We experiment on both the base (Olmo-7B-Base) and instruction-tuned (Olmo-7B-SFT) versions of the Olmo model Groeneveld et al. (2024). We also experiment on the Olmo-v1.7-Base model of the same size, which achieves better performance on a 4-way 0-shot formatted MCQA version of MMLU by making 0.01% of the pretraining data multiple-choice AI2 (2024). This performance gain is reported to occur at ~500B tokens of training Gu et al. (2024)— to investigate how model representations change once the formatted MCQA task is learned, we investigate one checkpoint before this point (350B training tokens), one checkpoint after (1T tokens), and the final checkpoint (~2T tokens).
Llama2.
We experiment with two sizes of base models from the Llama2 family Touvron et al. (2023): 7B and 13B, as well as the instruction-tuned RLHF-aligned 7B-Chat and 13B-Chat models.
All 7B (13B) models investigated have 32 (40) layers with 32 (40) attention heads per layer.
@@ 3.3 Identifying Consistent MCQA Models
We next isolate models capable of performing formatted MCQA. We design a set of 4 prompts to test whether models answer formatted MCQA prompts consistently, i.e., they are robust to the position and symbol of the correct answer choice.
We define consistency as always selecting one semantic answer for a dataset instance regardless of how the prompt for that instance is constructed. For example, for the two prompts ‘‘A:_B:_’’ and ‘‘A:_B:_’’, where bold denotes the location of the correct answer, the model is consistent if its two predictions are ‘‘A’’ followed by ‘‘B’’ (consistently correct) or ‘‘B’’ followed by ‘‘A’’ (consistently incorrect). Conversely, if the model always predicts the same token for both prompts, this indicates a preference towards that specific token or that specific position. We refer to this as being inconsistent (always ‘‘A’’ or always ‘‘B’’, depending on the model’s preference). Henceforth, since we are interested in consistent models, we will always average over ‘‘A:_B:_’’ and ‘‘A:_B:_’’ prompts when reporting accuracy, and refer to these as ‘‘A:_B:_’’ prompts.333Always predicting ‘‘A’’ or ‘‘B’’ will result in 50% accuracy. Lastly, to understand what behaviors we localize are specific to the letters used, we additionally include prompts ‘‘C:_D:_’’ and ‘‘C:_D:_’’.444When this prompt format is used, we format our context examples with labels ‘‘C’’, ‘‘D’’, ‘‘D’’.
We are interested in models that demonstrate learned ability to do the task of formatted MCQA with consistent symbol binding. We define such models as those that achieve at least 20% above random accuracy, i.e., average accuracy, on both ‘‘A:_B:_’’ and ‘‘C:_D:_’’. As Fig. 2 shows, only Olmo-7B-SFT and Llama2-13B-Chat achieve this criterion on all considered datasets. We focus on them in subsequent analysis.
Full results tables and statistics on consistency are included in Tables 1, 2, 3, 4 and 5 (Appendix), from which we make two observations: (I) Models that are consistently correct on ‘‘A:_B:_’’ also maintain this level of performance on ‘‘C:_D:_’’, despite the latter prompt being less likely to occur in the training data. (II) Models exhibit differing performance on the 3 datasets, indicating the datasets capture a meaningful range of difficulty.
@@ 4.1 Activation Patching
Activation patching, sometimes referred to as causal tracing Meng et al. (2022), causal mediation analysis Vig et al. (2020), or interchange intervention Geiger et al. (2020), is a method for performing mediation on neural network hidden states to localize which network components have a causal effect on a model’s prediction. The method involves performing the following steps:
@@
Run inference on a dataset instance that the model predicts correctly, say, one formatted as ‘‘A:_B:_’’ (). This will produce probabilities (high) and (low).
@@
Run inference on the same dataset instance, but in a different format, such as ‘‘A:_B:_’’, that the model still predicts correctly (). This will produce a second set of probabilities, (low) and (high). We store hidden state activations of interest.
@@
While running inference on , replace the output hidden state at the final token position and layer , , with the hidden state from the same layer and token position, but from the inference run of , . We measure and .
@@
We will use the notation to indicate patching a hidden state from into . Meng et al. (2022) proposed performing activation patching on the outputs of multi-layer perceptron (MLP) and multi-head self-attention (MHSA) functions; Todd et al. (2024) on the outputs of each weighted attention head.555MHSA output is a weighted sum of individual attention heads. See Section A.3 for background on Transformers. We also perform activation patching on these granularities (positions A and B in Fig. 10). Note two preliminary conditions for activation patching: the model predicts the correct answer for both and , and and do not have the same answer.
@@ 4.2 Vocabulary Projection
Residual connections have been shown to allow for the iterative refinement of features in neural networks Jastrzebski et al. (2018); Simoulin and Crabbé (2021). In addition to understanding which hidden states have the largest causal effect on predictions, it is useful to understand how predictions form in the vocabulary space defined by the dot product between the hidden state output by the LM, , and the unembedding matrix, . Recall that for an -layer autoregressive LM, model predictions are a probability distribution over each item in the vocabulary , , given by: where LN is layer normalization, and is the hidden state output by the final layer at the last token position.
Prior work (nostalgebraist, 2020; Geva et al., 2021, 2022a, 2022b; Dar et al., 2023; Katz and Belinkov, 2023; Yu et al., 2023; Merullo et al., 2024b, i.a.) has proposed to project any hidden state in the Transformer block of dimensionality to the vocabulary space, in order to inspect when and how models build up to via their internal representations. This technique, “vocabulary projection”, is equivalent to early exiting on the Transformer block at inference time (Dehghani et al., 2019; Elbayad et al., 2020; Schuster et al., 2021, 2022, i.a.). We plot the probability assigned to the two answer symbol tokens (‘‘A’’ and ‘‘B’’ or ‘‘C’’ and ‘‘D’’) as well as the maximum probability value in at various model states at final token position .

### v2

# Answer, Assemble, Ace: Understanding How LMs Answer Multiple Choice Questions

@@ Abstract
Multiple-choice question answering (MCQA) is a key competence of performant transformer language models that is tested by mainstream benchmarks. However, recent evidence shows that models can have quite a range of performance, particularly when the task format is diversified slightly (such as by shuffling answer choice order). In this work we ask: how do successful models perform formatted MCQA? We employ vocabulary projection and activation patching methods to localize key hidden states that encode relevant information for predicting the correct answer. We find that the prediction of a specific answer symbol is causally attributed to a few middle layers, and specifically their multi-head self-attention mechanisms. We show that subsequent layers increase the probability of the predicted answer symbol in vocabulary space, and that this probability increase is associated with a sparse set of attention heads with unique roles. We additionally uncover differences in how different models adjust to alternative symbols. Finally, we demonstrate that a synthetic task can disentangle sources of model error to pinpoint when a model has learned formatted MCQA, and show that logit differences between answer choice tokens continue to grow over the course of training.111Code is available at https://github.com/allenai/understanding_mcqa.
@@ 1 Introduction
Multiple-choice question answering (MCQA) is a mainstay of language model (LM) evaluation (Liang et al., 2022; Gao et al., 2023; Beeching et al., 2023), not in the least because it avoids the challenges of open-ended text evaluation. A prominent example is the MMLU benchmark (Hendrycks et al., 2021), which is considered a meaningful signal for new model releases (Lambert, 2024a; b).
While early LMs were evaluated on MCQA questions by treating the task as next-token-prediction and scoring each answer choice appended to the input independently, many benchmarks and datasets now evaluate models on their ability to produce the correct answer choice symbol in a single inference run when given the answer choices in the prompt (formatted MCQA; Figure 1 and Section 3.1). Indeed, it has been shown that capable models, and particularly those that have been instruction-tuned, perform well on formatted MCQA tasks (Liang et al., 2022; Robinson & Wingate, 2023; Wiegreffe et al., 2023). At the same time, despite their strong end-task performance, some models are not robust to seemingly inconsequential format changes, such as shuffling the position of the correct answer choice, changing letters that answers are mapped to, or using different symbols (Pezeshkpour & Hruschka, 2024; Alzahrani et al., 2024; Khatun & Brown, 2024).
@@
In this paper, we interpret internal components to uncover how models promote the prediction of specific answer choice symbols. We perform vocabulary projection and activation patching on three model families—Llama 3.1 (Dubey et al., 2024), Olmo 0724 (Groeneveld et al., 2024) and Qwen 2.5 (Yang et al., 2024)—and three datasets—MMLU (Hendrycks et al., 2021), HellaSwag (Zellers et al., 2019), and a copying task we create. Our findings, summarized in Figure 1, are:
@@
When models are correct, they both encode information needed to predict the correct answer symbol and promote answer symbols in the vocabulary space in a very similar fashion across tasks, even when their overall task performance varies.
@@
Answer symbol production is driven by a sparse portion of the network, namely, attention heads.
@@
The process for correctly answering more unnatural prompt formats is in some cases more complex: Olmo 7B Instruct and Qwen 2.5 1.5B models sometimes only begin producing the correct symbols for these prompts in later layers. We discover that the models’ hidden states initially assign high logit values to expected answer symbols (here, A/B/C/D) before switching to the symbols given in the prompt (here, random letters like Q/Z/R/X or O/E/B/P).
@@
A simple synthetic task can disentangle formatted MCQA performance from dataset-specific performance, and allows us to narrow down a point during training at which Olmo 7B learns formatted MCQA.
@@ 2 Related Work
Behavioral analyses of LM abilities to answer formatted MCQA questions prompt models in a black-box manner by constructing different input prompts and observing how they affect models’ predictions (Wiegreffe et al., 2023; Pezeshkpour & Hruschka, 2024; Sun et al., 2024; Zheng et al., 2024; Alzahrani et al., 2024; Khatun & Brown, 2024; Balepur et al., 2024; Wang et al., 2024). These methods and findings are complementary to ours.
Efforts to interpret transformer LMs mechanistically222In the narrow technical sense of the term (Saphra & Wiegreffe, 2024). for MCQA are limited. Most similar to ours is Lieberum et al. (2023), who study what attention heads at the final token position are doing, and isolate a subset they coin “correct letter heads”. They demonstrate how these heads attend to answer symbol representations at earlier token positions and promote the correct symbol based on its position in the symbol order, though they show this behavior does not hold entirely for answer choices other than A/B/C/D. Their experiments are limited to one task (MMLU) and one closed-source model (Chinchilla-70B). Apart from broadening the scope of analysis substantially, we also analyze model behavior over the course of training and when it is poor, and create a synthetic task to disentangle task-relevant knowledge from the ability to answer formatted MCQA questions. We additionally put forward a hypothesis for how models adapt when other answer choice symbols are used.
Li & Gao (2024) use vocabulary projection and study the norms of weighted value vectors in multi-layer perceptrons of GPT-2 models in order to localize the production of A, which GPT-2 models are biased towards producing when prompted 0-shot. They find that updating a single value vector can substantially reduce this label bias. Unlike that work, we investigate how models do symbol binding when they are performant at formatted MCQA; label bias is one reason why a model may not be correctly performing symbol binding.
Outside of MCQA, many works investigate how models internally build up representations that lead to their predictions, such as on factual recall (Geva et al., 2021; Dai et al., 2022; Geva et al., 2022b; Meng et al., 2022; Geva et al., 2023; Yu et al., 2023; Merullo et al., 2024a) or basic linguistic (Vig et al., 2020; Wang et al., 2023; Merullo et al., 2024b; Prakash et al., 2024; Todd et al., 2024) or arithmetic (Stolfo et al., 2023; Hanna et al., 2023; Wu et al., 2023) tasks. MCQA differs because it is both directly represented in real benchmarks and involves multiple reasoning steps.
@@ 3 Formatted MCQA
3.1 Task Notation
Formatted MCQA is a prompt format in which possible answer choices are presented to the model as an enumerated list, each associated with an answer choice symbol. All of our dataset instances are either 0-shot or 3-shot with 4 answer choices, formatted as follows:
For each of the following phrases, select the best completion. <optional in-context examples of the same format> Phrase: Corn is yellow. What color is corn? Choices: A. yellow B. grey C. blue D. pink The correct answer is:
where A, B, C, or D (A in this example).333The actual answer tokens include preceding spaces in some of the tokenizers of the models we study. Models parameterized by are evaluated on their ability to assign higher probability to the symbol associated with the correct answer choice (here, A) as opposed to the alternative symbols (here, B, C, and D). We compute . To perform well at this task, models must not only predict the correct answer phrase (yellow in the above example), but then map it to its corresponding symbol (A). Robinson & Wingate (2023) refers to this capability as symbol binding.
@@ 3.2 Datasets
We experiment on two challenging real-world multiple-choice datasets from LLM benchmarks: HellaSwag (Zellers et al., 2019) and MMLU (Hendrycks et al., 2021). Details are given in Section A.1. We additionally use a prototypical colors dataset (Norlund et al., 2021) to create a synthetic 4-way task disentangling dataset-specific knowledge from the ability to perform symbol binding: Copying Colors from Context (‘Colors’). The dataset consists of instances such as Corn is, yellow. We include in the context, so the model’s only task is to produce the symbol associated with that answer choice given three distractors randomly selected from the ten colors present in the dataset (i.e., perform symbol binding). We use 3 instances as in-context examples and the remaining 105 as our test set. An example formatted instance is given above.
We validate that all models achieve 100% 3-shot accuracy on a generative version of the Colors task: we provide instances in the format ‘‘ Corn is yellow. What color is corn?’’ and take the first greedy-decoded token as the prediction (i.e., we expect the model to produce ‘‘yellow’’).
@@ 3.3 Models
Olmo family.
We experiment on the base (Olmo 0724 7B) and instruction-tuned (Olmo 0724 7B Instruct) versions of the most recent (0724) release of the Olmo model (Groeneveld et al., 2024). These have 32 layers with 32 attention heads per layer. We additionally benchmark performance on the smaller Olmo 0724 1B and supervised finetuned Olmo 0724 7B SFT models.
Llama family.
We experiment with the smallest Llama 3.1 models: Llama 3.1 8B base model and Llama 3.1 8B Instruct (Dubey et al., 2024). These have 32 layers with 32 attention heads per layer.
Qwen family.
To study smaller performant models, we include the base and instruct versions of the 0.5B and 1.5B Qwen 2.5 models (Yang et al., 2024). The 1.5B model has 28 layers with 12 attention heads per layer.
@@ 3.4 Identifying Consistent MCQA Models
We next isolate models capable of performing formatted MCQA by testing whether models answer formatted MCQA prompts consistently, i.e., they are robust to the position and symbol of the correct answer choice. For each dataset instance, we construct four versions where we vary the location of the correct answer string, and thus . We denote these as A/B/C/D, A/B/C/D, A/B/C/D, and A/B/C/D, where bold indicates the location of the correct answer. We average over these four prompts when reporting accuracy, and refer to this collective set of instances of size as A/B/C/D prompts.444Always predicting a single letter results in 25% accuracy when averaged across correct answer positions. This ensures that models with a strong bias towards a particular symbol are penalized. Lastly, to understand what behaviors we localize are specific to the letters used, we additionally include prompts Q/Z/R/X and 1/2/3/4.555When this prompt format is used, we format the in-context examples with the same symbols. We evaluate each model per dataset instance on this set of 12 total prompts.
Results are in Fig. 2, where the minimum performance across A/B/C/D, Q/Z/R/X, and 1/2/3/4 is plotted. Models exhibit differing performance on the 3 datasets, indicating the datasets capture a meaningful range of difficulty. We additionally observe that many models achieve perfect or near-perfect performance on Colors while being far from it on the other two datasets, indicating that error can be attributed to dataset difficulty as opposed to symbol binding. Finally, most (but not all) models are capable of performing symbol binding on the Colors dataset with all 3 sets of answer choice symbols, despite the Q/Z/R/X prompt being far less likely to occur in the training data.666For models that do not perform well on the Colors task, this is due to the MCQA format, since all models can copy and produce the color string from the context with 100% accuracy (Section 3.2).
We select the best-performing model from each model family for subsequent analysis: Olmo 0724 7B Instruct, Llama 3.1 8B Instruct, and Qwen 2.5 1.5B. They perform all 3 tasks sufficiently above random accuracy (20% above random, when averaged across prompts and answer positions).
@@ 4.1 Activation Patching
Activation patching, sometimes referred to as causal tracing (Meng et al., 2022), causal mediation analysis (Vig et al., 2020), or interchange intervention (Geiger et al., 2020), is a method for performing mediation on neural network hidden states to localize which network components have a causal effect on a model’s prediction. The method involves performing the following steps:
@@
Run inference on a dataset instance that the model predicts correctly, say, one formatted as A/B/C/D (). This will produce scores (higher) and , , (lower).
@@
Run inference on the same dataset instance, but in a different format, such as A/B/C/D, that the model still predicts correctly (). This will produce a second set of scores (higher) and , , (lower). We store hidden state activations of interest.
@@
While running inference on , replace the output hidden state at the final token position and layer , , with the hidden state from the same layer and token position, but from the inference run of , . We measure and plot , , , and .
@@
We use the notation to indicate patching a hidden state from into . Meng et al. (2022) proposed performing activation patching not only on layerwise outputs, but also on the outputs of multi-layer perceptron (MLP) and multi-head self-attention (MHSA) functions; Todd et al. (2024) extend the latter by patching the outputs of each weighted attention head.777MHSA output is a weighted sum of individual attention heads. See Section A.2 for background on the Transformer architecture. We do the same. Note two preliminary conditions for activation patching: the model predicts the correct answer for both and , and and do not have the same answer. The score can be either a logit or probit value; we discuss this in Section 4.3.
@@ 4.2 Vocabulary Projection
Residual connections have been shown to allow for the iterative refinement of features in neural networks (Jastrzebski et al., 2018; Simoulin & Crabbé, 2021). In addition to understanding which hidden states have the largest causal effect on predictions, it is useful to understand how predictions form in the vocabulary space defined by the dot product between the hidden state output by the LM, , and the unembedding matrix, . Recall that for an -layer autoregressive LM, model predictions are derived from logit values assigned to each item in the vocabulary , , given by: where LN is layer normalization, and is the hidden state output by the final layer at the last token position. The scores are then further normalized into probits that sum to 1 over the vocabulary: .
Prior work (nostalgebraist, 2020; Geva et al., 2021; 2022a; 2022b; Dar et al., 2023; Katz & Belinkov, 2023; Yu et al., 2023; Merullo et al., 2024a, i.a.) has proposed projecting any hidden state in the Transformer block of dimensionality to the vocabulary space, in order to inspect when and how models build up to and/or via their internal representations. This technique, “vocabulary projection”, also sometimes called “logit lens” when logits are used, is equivalent to early exiting on the Transformer block at inference time (Dehghani et al., 2019; Elbayad et al., 2020; Schuster et al., 2021; 2022, i.a.). We plot the values assigned to the answer symbol tokens, as well as the maximum values in and assigned to any other token, at model states from each layer at the final token position . The “max other token” line delineates when scores assigned to the tokens of interest reach the top of the vocabulary ranking. The “logit difference” line plots the gap between the predicted token and the next-largest of the answer choices (akin to the scoring rule described in Section 3.1).
4.3 Logits vs. Probits
Logits are a useful choice of metric for activation patching and vocabulary projection, in part because they are capable of detecting model components which demote particular answer choices by assigning lower-than-average logit values (Zhang & Nanda, 2024); these can become indistinguishable from average logit values (i.e., both on or near 0) after Softmax normalization. They also measure the direct additive contribution that each model component makes to the residual stream, and thus have a loosely causal interpretation, assuming no layer normalization on model outputs (Lieberum et al., 2023; Zhu et al., 2024). On the other hand, it is valuable to measure scores assigned to the tokens of interest (such as answer symbol tokens) with respect to scores assigned to other tokens in the vocabulary and on a normalized 0-1 scale, as probits do. This 1) gives us insight into the rank of the tokens of interest in vocabulary space, which affects greedy generation; 2) allows us to compare processing patterns across different models, which often have different absolute logit scales, and 3) allows us to disentangle logit changes that are specific to the tokens of interest as opposed to general trends to the entire model vocabulary. We thus use both metrics because they provide complementary information.